Show that transformation 𝑇:𝑅

2 → 𝑅

2 defined by 

𝑇(𝑥, 𝑦) = (𝑎𝑥 + 𝑏𝑦, 𝑐𝑥 + 𝑑𝑦),


1
Expert's answer
2021-12-14T03:27:08-0500

Given

"\begin{aligned} T: R^{2} \rightarrow R^{2} & \text { such map } \\ & T(x, y)=(a x+b y,(x+d y) .\end{aligned}"

"T(x, y)=(a x+b y, c x+d y) \cdot u, v \in R^{2}\ \text{such that}"

 "\begin{aligned}

T(x, y) &=(u, v) \cdot \\

\text { i.e } \quad a x+b y &=u \\

c x+d y &=v \\

\left[\begin{array}{ll}

a & b \\

c & d

\end{array}\right]\left[\begin{array}{l}

x \\

y

\end{array}\right] &=\left[\begin{array}{l}

u \\

v

\end{array}\right] \\

\left[\begin{array}{l}

x \\

y

\end{array}\right] &=\left[\begin{array}{ll}

a & b \\

c & d

\end{array}\right]^{-1}\left[\begin{array}{l}

u \\

v

\end{array}\right] . \\

\left[\begin{array}{l}

x \\

y

\end{array}\right] &=\left[\begin{array}{ll}

d & -b \\

-c & a

\end{array}\right] \frac{1}{a d-b c}\left[\begin{array}{l}

u \\

v

\end{array}\right]

\end{aligned}"

There exists (x, y) such that

 "T(x, y)=(u, v) \text { iff } a d-b c \neq 0 \text {. }"

a, b, c, d are scalars such that "a d-b c \neq 0."